OpenStudy (anonymous):

what are the roots of the equation x^2-5x+6+0

OpenStudy (anonymous):

first you have to factor the equation which is (x-2)(x-3) = 0 then use the zero product property to solve for the factors x-2 = 0 x-3 = 0 so your answers are x=2 x=3